Given the quadratic equation x^2-15=2k(x-4), find k if there are two real and equal roots.
Found 2 solutions by stanbon, bucky:
Answer by stanbon(75887)

Given the quadratic equation x^2-15=2k(x-4), find k if there are two real and equal roots.
-----------
The discriminant will be zero.
-------------
x^2-15=2kx-8k
x^2-2kx+8k-15=0
---------------
Discriminant: b^2-4ac
(2k)^2-4*1(8k-15) = 0
4k^2-32k+60 =0
k^2-8k+15 = 0
(k-3)(k-5)=0
k = 3 or k=5
================
Cheers,
Stan H.

Answer by bucky(2189)

Clarification to Stanbon answer:
.
The last step results in two answers. K = 0 is not correct. If you look closely at the Stanbon
answers, the results should be K = +3 and K = +5. Just a typo.
.
The methodology shown is the way to do the problem ...
.